Josh is skating a pond. He moves at 10 mph going from one end to the other, but on his way back he comes 5 mph. It took him one hour. How long is the pond?
i tried 10(x-1) + 5(x)

Answer by Alan3354(69443)

Josh is skating a pond. He moves at 10 mph going from one end to the other, but on his way back he comes 5 mph. It took him one hour. How long is the pond?
--------------
The avg speed for the round trip is
2*10*5/(10+5) = 100/15 mi/hr
d = r*t = (100/15) mi/hr * 1hr = 20/3 miles RT
--> 10/3 miles one-way
-------------
Avg speed for a round trip is alway 2*r1*r2/(r1 + r2)
You can work that out, or take my word for it.